Working alone, it takes Sue 13 hours to pick 40 bushels of apples. Kim can pick the same amount in 15 hours. If they work together, how long would it take them?
Thamk you!

Answer by richwmiller(17219)

x/13+x/15=1
15x+13x=13*15
28x=195
x=195/28